The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

           mcnews@astro.uiuc.edu
http://www.astro.uiuc.edu/projects/mcnews/MCNews.html
   http://www.astro.uni-bonn.de/~mcnews/

No. 48 October 24, 2000

Contents

News  1
Abstracts of 14 refereed papers  2
Abstracts of 4 non-refereed papers 11
2 job opportunities  13
1 meeting announcement  15

News

OGLE BVI Maps of the LMC available over the Internet

The Optical Gravitational Lensing Experiment (OGLE) team announces the release of “OGLE BV I maps of Dense Stellar Regions – The Large Magellanic Cloud.” The maps contain BV I photometric and astrometric data of about 7 million stars from the central regions of the LMC (about 5.6 square degrees) observed during the OGLE microlensing survey.

The maps can be retrieved from the OGLE Internet archive:

 http://www.astrouw.edu.pl/~ogle
ftp://sirius.astrouw.edu.pl/ogle/ogle2/maps/lmc/

or its US mirror

http://bulge.princeton.edu/~ogle
ftp://bulge.princeton.edu/ogle/ogle2/maps/lmc/

Andrzej Udalski for the OGLE team.

Abstracts of Refereed Papers

The critical velocity effect as a cause for the Hα emission from the Magellanic Stream

C. Konz^{1}, H. Lesch^{1}, G.T. Birk^{1}, and H. Wiechen^{1}

1 Centre for Interdisciplinary Plasma Science (CIPS), Institut für Astronomie und Astrophysik der Universität München, Scheinerstraße 1, D-81679 München, Germany

Observations show significant Hα-emissions in the Galactic halo near the edges of cold gas clouds of the Magellanic Stream. The source for the ionization of the cold gas is still a widely open question. In our paper we discuss the critical velocity effect as a possible explanation for the observed Hα-emission. The critical velocity effect can yield a fast ionization of cold gas if this neutral gas passes through a magnetized plasma under suitable conditions. We show that for parameters that are typical for the Magellanic Stream the critical velocity effect has to be considered as a possible ionization source of high relevance.

Accepted by: The Astrophysical Journal appear

For preprints, contact konz@usm.uni-muenchen.de
Also available from the URL http://www.arXiv.org/abs/astro-ph/0010164

The Optical Gravitational Lensing Experiment.

BVI Maps of Dense Stellar Regions. II. The Large Magellanic Cloud

A. Udalski^{1}, M. Szymański^{1}, M. Kubiak^{1}, G. Pietrzyński^{1,2}, I. Soszyński^{1}, P. Woźniak^{3}, and K. Żebruń^{1}

1 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland 2 Universidad de Concepción, Departamento de Fisica, Casilla 160–C, Concepción, Chile 3 Princeton University Observatory, Princeton, NJ 08544-1001, USA

We present the BVI photometric maps of the Large Magellanic Cloud. They contain BVI photometry and astrometry of more than 7 million stars from the central parts of the LMC. The data were collected during the second phase of the OGLE microlensing project. We discuss the accuracy of the data and present color-magnitude diagrams of all 26 fields observed by OGLE in the LMC.

The BVI maps of the LMC are accessible electronically for the astronomical community from the OGLE Internet archive.

Accepted by: Acta Astronomica Vol. 50, 307.

For preprints, contact udalski@astrouw.edu.pl
Also available from the URL http://www.arXiv.org/abs/astro-ph/0010150

A Search for Wolf-Rayet Stars in the Small Magellanic Cloud

Philip Massey^{1} and Alaine S. Duffy^{2}

1 Lowell Observatory, 1400 W. Mars Hill Road, Flagstaff, AZ 86001, USA 2 Gettysburg College, Box 756, Gettysburg, PA 17325, USA

We report on a comprehensive search for Wolf-Rayet (W-R) stars in the SMC using interferencefilter imaging. Photometry of over 1.6 million stellar images on multiple, overlapping fields covering 9.6 square-degrees found the previously known W-Rs at very high significance levels, two known Oftype stars, plus additional candidates which we examined with slit spectroscopy. We discovered two new Wolf-Rayet stars, both of type “WN3+abs”, bringing the total number in the SMC to 11. We discuss their spectra, as well as reclassifying the previously known ones with our new data. Our survey also revealed 4 newly found Of-type stars, including one of O5e If+ type, one of the earliest-type stars known in the SMC. Another newly identified Of star is AV 398 (O8 If), a star often used in extinction studies under the assumption that it is of early B type . We recover S18 (AV154), a B[e] star whose spectrum currently lacks He II λ4686 emission, but which must have had strong emission a year earlier; we compare this star to S Dor, suggesting it is indeed a luminous blue variable (LBV). We also find a previously unknown symbiotic star whose spectrum is nearly identical to the Galactic symbiotic AG Dra. More important, perhaps, than any of these discoveries is the demonstration that there is not a significant number of W-Rs waiting to be discovered in the SMC. The number of W-Rs is a factor of 3 times lower in the SMC (per unit luminosity) than in the LMC. This strongly suggests in that at the low metallicity which characterizes the SMC that only the most massive stars can evolve to W-R type.

Submitted to: Astrophysical Journal For preprints, contact massey@lowell.edu

Also available from the URL http://www.arXiv.org/abs/astro-ph/0010420
                          ftp.lowell.edu/pub/massey/smcwr.ps.gz

HST/STIS Observations of the Wolf-Rayet Star HD 5980 in the Small Magellanic Cloud: II. The Interstellar Medium Components

G. Koenigsberger^{1}, L. Georgiev^{1}, M. Peimbert^{1}, N. R. Walborn^{2}, R. Barbá^{3}, V.S.

Niemela^{3}, N. Morrell^{3},Z. Tsvetanov^{4}, and R. Schulte-Ladbeck^{5}

1 Instituto de Astronomía, UNAM, Apdo. Postal 70-264, Mexico D.F. 04510 2 Space Telescope Science Institute, 3700 San Martin Dr., Baltimore, MD 21218, USA 3 Facultad de Ciencias Astronómicas y Geofísicas, Universidad Nacional de la Plata, Paseo del Bosque S/N, B19000FWA La Plata, Argentina 4 Johns Hopkins University, Baltimore, MD 21218, USA 5 Department of Physics and Astronomy, University of Pittsburgh, Pittsburgh, PA 15260, USA

Observations of the interstellar and circumstellar absorption components obtained with the Hubble Space Telescope Imaging Spectrograph (STIS) along the line-of-sight toward the WR/LBV system HD 5980 in the Small Magellanic Cloud are analyzed. Velocity components from C I, C I*, C II, C II*, C IV, N I, N V, O I, Mg II, Al II, Si II, Si II*, Si III, Si IV, S II, S III, Fe II, Ni II, Be I, Cl I, and CO are identified, and column densities estimated. The principal velocity systems in our data are: 1) ISM components in the Galactic disk and halo (it V_{helio}= 1.1±3, 9±2 km s^{−1}); 2) ISM components in the SMC (v_{helio} =+87±6, +110±6, +132±6, +158±8, +203±15 km s^{−1}); 3) SMC supernova remnant SNR0057-7226 components (v_{helio}=+312±3, +343±3, +33, +64 km s^{−1}); 4) circumstellar (CS) velocity systems (v_{helio} =−1020,−840,−630,−530,−300 km s^{−1}); and 5) a possible system at −53±5 km s^{−1} (seen only in some of the Si II lines and marginally in Fe II) of uncertain origin. The supernova remnant SNR0057-7226 has a systemic velocity of +188 km s^{−1}, suggesting that its progenitor was a member of the NGC 346 cluster. Our data allow estimates to be made of T_{e}∼40,000 K, n_{e}∼100 cm^{−3}, N(H)∼(4-12)×10^{18} cm^{−2} and a total mass between 400 and 1000 M_{⊙} for the SNR shell. We detect C I absorption lines primarily in the +132 and +158 km s^{−1} SMC velocity systems. As a result of the LBV-type eruptions in HD 5980, a fast-wind/slow-wind circumstellar interaction region has appeared, constituting the earliest formation stages of a wind-blown H II bubble surrounding this system. Variations over a timescale of one year in this CSM structure are detected.

Accepted by: The Astronomical Journal

For preprints, contact gloria@astroscu.unam.mx

The X-ray Remnant of SN1987A

David N. Burrows^{1}, Eli Michael^{2}, Una Hwang^{3}, Richard McCray ^{2}, Roger A. Chevalier^{4}, Robert Petre^{3}, Gordon P. Garmire^{1}, Stephen S. Holt^{3}, John A. Nousek^{1}

1 Department of Astronomy & Astrophysics, Penn State University, University Park, PA 16802, USA 2 JILA, University of Colorado, Boulder, CO 80309-0440, USA 3 NASA/Goddard Space Flight Center, Greenbelt, MD 20771, USA 4 Dept. of Astronomy, University of Virginia, P.O. Box 3818, Charlottesville, VA 22903, USA

We present high resolution Chandra observations of the remnant of SN1987A in the Large Magellanic Cloud. The high angular resolution of the Chandra X-ray Observatory (CXO) permits us to resolve the X-ray remnant. We find that the remnant is shell-like in morphology, with X-ray peaks associated with some of the optical hot spots seen in HST images. The X-ray light curve has departed from the linear flux increase observed by ROSAT, with a 0.5-2.0 keV luminosity of 1.5 × 10^{35} erg/s in January 2000. We set an upper limit of 2.3 × 10^{34} ergs/s on the luminosity of any embedded central source (0.5 - 2 keV). We also present a high resolution spectrum, showing that the X-ray emission is thermal in origin and is dominated by highly ionized species of O, Ne, Mg, and Si.

Accepted by: The Astrophysical Journal (Letters)

For preprints, contact burrows@astro.psu.edu
Also available from the URL http://www.arXiv.org/abs/astro-ph/009265
                          ftp.astro.psu.edu://pub/Publications/burrows/SN1987a.pdf

ASCA Observations of the Twin Supernova Remnants in the Large Magellanic Cloud, DEM L316

Mamiko Nishiuchi^{1}, Jun Yokogawa^{1}, Katsuji Koyama^{1} and John P. Hughes^{2}

1 Department of Physics, Graduate School of Science, Kyoto University, Sakyo-ku, Kyoto, 606-8502, Japan 2 Department of Physics and Astronomy, Rutgers University, 136 FrelinghuysenRoad, Piscataway, NJ 08854-8019 U.S.A.

We report results from an ASCA X-ray study of DEM L316, an emission nebula in the Large Magellanic Cloud (LMC) consisting of two closely-spaced supernova remnants (SNRs). The SIS image shows separate X-ray sources located at the positions of the two radio- and optically-emitting SNR shells, 0547−69.7A and 0547−69.7B (hereafter, shell A and B). The individual X-ray spectrum of each shell is well described by optically-thin thermal emission, although the characteristics of the emission differ in important details between them. Shell A exhibits strong iron L emission, which we attribute to the presence of iron-rich ejecta leading to the suggestion that this SNR originates from a Type Ia SN, an explosion of a moderate mass progenitor. Shell B, on the other hand, has a chemical composition similar to that of the interstellar medium of the LMC and so its X-ray emission is dominated by swept-up matter. The different spectral features strongly disfavor the hypothesis that the two shells are due to one supernova explosion into an interconnected bubble. We could not obtain the evidence which positively supports the collision between two SNRs.

Accepted by: Publications of the Astronomical Society of Japan

For preprints, contact mamiko@cr.scphys.kyoto-u.ac.jp
Also available from the URL http://www.arXiv.org/abs/astro-ph/009318
                          http://www-cr.scphys.kyoto-u.ac.jp/MC/

Morphology and Evolution of the LMC Planetary Nebulae

Richard A. Shaw^{1}, Letizia Stanghellini^{1}, Max Mutchler^{1}, Bruce Balick^{2} and J. Chris

Blades^{1}

1 Space Telescope Science Institute, Baltimore, MD 21218, USA 2 University of Washington, Seattle, WA 98195, USA

The LMC is ideal for studying the co-evolution of planetary nebulae (PNe) and their central stars, in that the debilitating uncertainties of the Galactic PN distance scale and selection biases from attenuation by interstellar dust do not apply. We present images and analyze slit-less spectra which were obtained in a survey of Large Magellanic Cloud PNe. These data on 29 targets were obtained with HST using the Space Telescope Imaging Spectrograph. The data permit us to determine the nebular dimensions and morphology in the monochromatic light of several emission lines, including those that have traditionally been used for morphological studies in the Galaxy: Hα, [N ii] 6583 Å and [O iii] 5007 Å, plus others of varying ionization, including [O i], He i, and [S ii]. Together with the 31 resolved LMC PNe for which monochromatic images exist in the HST archive, these data show that the incidence of non-symmetric nebulae, including bipolar nebulae (which is an indicator of Population I ancestry in the Galaxy), is significantly higher than that reported for the Galaxy. The onset of asymmetric features appears even in very young nebulae (with dynamical ages of ∼ 1400 yr), suggesting that at least the gross features of the nebular morphology may be more closely tied to PN formation, and that subsequent shaping of the expanding envelope by the radiation field and wind from the central star may play the lesser role of amplifying these gross features. There is some evidence of evolution between two morphological types, in the sense that bipolar core (BC) nebulae may evolve to pure bipolars late in the PN lifetime.

Accepted by: The Astrophysical Journal For preprints, contact shaw@stsci.edu

Also available from the URL http://xxx.lanl.gov/ps/astro-ph/0010200

LMC and SMC Cepheids: combining ogle with denis and 2mass

M.A.T. Groenewegen^{1,2}

1 Max-Planck Institut für Astrophysik, Karl-Schwarzschild-Straße 1, D-85740 Garching, Germany 2 Currently at: ESO (European Southern Observatory), EIS-team, Karl-Schwarzschild-Straße 2, D-85748 Garching, Germany

I cross-correlate the ogle database of Cepheids in the Large and Small Magellanic Clouds (MCs) with the second incremental release of the 2mass JHK_{s} survey, and the denis IJK_{s} point source catalog towards the MCs.

Of the 3384 Cepheids in ogle in both Clouds, 1745 have a counterpart in the 2mass survey within an 1′′ radius and good-photometry in all three bands. Only 173 have a counterpart in the denis survey within an 1′′ radius and good-photometry in all three bands. The reason for this difference is that the limiting magnitudes of the 2mass survey are considerably fainter than for the denis survey.

The standard stars of the Carter IR photometric system are also correlated with the 2mass survey to derive transformation equations from the natural 2mass system to the Carter-system.

In order to describe to first order the geometrical configuration of the MCs a plane is fitted to the reddening-free Wesenheit index (W = I − 1.55 × (V − I)) and the inclination and position angle of the line-of-nodes are determined. For the LMC an inclination angle of 18 ± 3^{∘} is derived, somewhat smaller that previous determinations. For the SMC a value of i = 68 ± 2^{∘} is derived in good agreement with previous determinations. These results are being used to take out the effect that some Cepheids are closer to the observer than others. For the LMC this effect is small over the area covered by the ogle survey, but for the SMC P L-relations with a reduced scatter are derived in this way. Nevertheless, the dispersion in the P L-relation in W for the SMC is larger than that for the LMC, indicating a significant intrinsic depth of the SMC of about 14 kpc (front-to-back). The depth of the LMC is found to be small compared to the intrinsic scatter in the P L-relation.

Using single-epoch 2mass data for the ogle Cepheids P L-relations are fitted in JHK (on the Carter system) for both Fundamental mode (FU) and First-overtone (FO) pulsators, taking out the effect of the inclined disk. Because of the large number of stars available, this results in the most accurately determined slopes in the infrared up to now, and slopes for FO pulsators derived for the first time. This provides additional constraints for theoretical models.

The 16 derived P L-relations (in W JHK for FU and FO pulsators for LMC and SMC) are used to derive a relative distance modulus Δ_{SMC−LMC} = 0.50 ± 0.02 assuming no metallicity correction. Two recent theoretical models predict different metallicity corrections, indicating a systematic uncertainty of up to 0.1 mag.

Using the database of Galactic Cepheids observed with hipparcos and the method outlined in Groenewegen & Oudmaijer (2000), the zero points of Galactic P L-relations for FU pulsators in W and K are derived, using the newly derived slopes for LMC and SMC pulsators.

Combining the zero points of the Galactic and MCs P L-relations distance moduli of 18.60 ± 0.11 (based on W) and 18.55 ± 0.17 (based on K) to the LMC, and of 19.11 ± 0.11 (W) and 19.04 ± 0.17 (K) to the SMC are derived (without taking into account possible metallicity corrections).

Accepted by: Astronomy & Astrophysics For preprints, contact mgroenew@eso.org

Also available from the URL http://xxx.lanl.gov/ps/astro-ph/0010298
                          http://www.hq.eso.org/^{∼}mgroenew/

S Doradus variables in the Galaxy and the Magellanic Clouds

A.M. van Genderen^{1}

1 Leiden Observatory, Postbus 9513, NL-2300RA Leiden, The Netherlands

The goal in writing this paper is five fold: (1) to summarize the scientific achievements in the 20th century on S Dor variables (or LBVs); (2) to present an inventory of these variables in the Galaxy and the Magellanic Clouds with a description of their physical state and instability properties; (3) to emphasize the photometric achievements of the various types of instabilities. Generally this seems to be a neglected item resulting in a number of misunderstandings continuously wandering through literature; (4) to investigate the structure of the S Dor-area on the HR-diagram; (5) to estimate the total numbers of S Dor variables in the three stellar systems.

The position of the strong-active S Dor variables in minimum brightness obey the following linear relation on the HR-diagram:

logL/L_{⊙} = 1.37 log T_{eff} − 0.03

The relatively small dispersion of less active and supposed ex- and dormant S Dor variables with respect to this relation is twice as large at the blue side than at the red side. This might be caused by evolution to the WR stage and/or to high rotation.

S Dor variables can be subject to five types of instabilities: the very rare genuine eruptive episodes (the ‘SD-eruptions’), two different brightening phases caused by slow pulsations (the ‘SD-phases’): one on a time scale of years, the other on a time scale of decades at a more or less constant luminosity and two types of microvariations: one on a time scale of weeks, the other on a time scale of about 100 d. So far, no periodicities of light curve characteristics of any of these instabilities have ever been found. The durations of active and non-active stages are estimated for about half of the sample based on scattered magnitude estimations such as from historical records, and on monitoring campaigns. It would be a misunderstanding to believe that all S Dor variables should be always spectacular. It is estimated that most of them will be not spectacular at all for at least 70% of their lifetime as an S Dor variable.

Accepted by: Astronomy & Astrophysics Supplements

For preprints, contact genderen@strw.leidenuniv.nl

ROSAT HRI catalogue of X-ray sources in the SMC region

Sasaki M., Haberl F., Pietsch W.

Max-Planck-Institut für extraterrestrische Physik, Giessenbachstraße, Postfach 1312, 85741 Garching, Germany

During the operational phase of the ROSAT satellite between 1990 and 1998 the X-ray telescope pointed 71 times to the Small Magellanic Cloud (SMC) for observations with the High Resolution Imager (HRI), covering a field of 5^{∘} × 5^{∘}. From this data a catalogue of 121 discrete X-ray sources was derived. By cross-correlating the source catalogue with the SIMBAD data base and the TYCHO catalogue, the systematic positional error of the HRI source positions could be reduced. In total the X-ray position for 99 HRI sources was corrected yielding positional errors between 1′′ and 16′′. The HRI catalogue was also cross-correlated with the catalogues derived from the ROSAT Position Sensitive Proportional Counter (PSPC) pointings (Kahabka et al. 1999; Haberl et al. 2000). For 75 HRI sources PSPC counterparts were found and thus hardness ratios are given. With the help of the information obtained from the cross-correlations 56 HRI sources were identified with objects of known or proposed nature. Four foreground stars, six supernova remnants, four supersoft sources, 12 X-ray binaries, and one AGN were detected by the HRI. Based on the existence of a likely optical counterpart or properties like hardness ratio and X-ray to optical flux ratio, further 15 HRI sources were classified into different source types.

Accepted by: Astronomy & Astrophysics Supplements For preprints, contact manami@mpe.mpg.de

Also available from the URL http://www.arXiv.org/abs/astro-ph/009390

The Optical Gravitational Lensing Experiment. Ages of about 600 Star Clusters from the LMC

G. Pietrzyński^{1,2} and A. Udalski^{1}

1 Universidad de Concepción, Departamento de Fisica, Casilla 160–C, Concepción, Chile 2 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland

We present results of age determination based on the standard procedure of isochrone fitting for about 600 star clusters younger than about 1.2 Gyr from the central parts of the LMC. Comparison of age distributions of star clusters from the LMC and the SMC shows that cluster formation histories are different in these galaxies. The age distribution of the LMC clusters reveals bursty nature of cluster formation in this galaxy, with contrast to the relatively uniform distribution of cluster ages in the SMC. We detected three extended peaks in the distribution of ages of LMC clusters, located at about 7 Myr, 125 Myr and 0.8 Gyr. All detected peaks have complex structure. While the structure of the youngest and the oldest peaks may be spurious due to accuracy of age determination, in the middle peak two evident sub-peaks at 100 Myr and 160 Myr are clearly seen. Similar peaks are seen in the distribution of ages of clusters from the SMC, which indicates that increased cluster formation ratio during these periods might be caused by the last encounter between the Magellanic Clouds. The sample of clusters from the LMC is overabundant in very young (< 10 Myr) and old (> 400 Myr) clusters, and underabundant in clusters having ages in the range from 80 to 15 Myr with respect to the studied sample of clusters from the SMC.

Accepted by: Acta Astronomica Vol. 50, 337.

For preprints, contact pietrzyn@astrouw.edu.pl
Also available from the URL http://www.arXiv.org/abs/astro-ph/0010360

The Optical Gravitational Lensing Experiment. Multiple Cluster Candidates in the Large Magellanic Cloud

G. Pietrzyński^{1,2} and A. Udalski^{1}

1 Universidad de Concepción, Departamento de Fisica, Casilla 160–C, Concepción, Chile 2 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland

We present 100 multiple cluster candidates selected from the OGLE catalog of star clusters in the Large Magellanic Cloud. Statistical analysis shows that the significant fraction of these objects may constitute physical systems. Coeval ages of 102 components of multiple objects suggest their common origin. 53 components have very different ages. The comparison of the population of multiple clusters candidates from the SMC and the LMC shows that: a) distributions of sizes and ages of multiple and single clusters from the Magellanic Clouds are very similar, b) the difference of sizes of components of a given system is small, c) in both distributions of separation of multiple clusters from the LMC and SMC two peaks are seen at about 9 pc and 15 pc, d) both age distributions reveal peaks around 100 Myr, which may be connected with the last encounter of the LMC and the SMC.

Accepted by: Acta Astronomica Vol. 50, 355.

For preprints, contact pietrzyn@astrouw.edu.pl
Also available from the URL http://www.arXiv.org/abs/astro-ph/0010294

Age-metallicity relation and Chemical evolution of the LMC from UVES spectra of Globular Cluster giants

V. Hill^{1} P. François^{1} M. Spite^{2} F. Primas^{1} and F. Spite^{2}

1 European Southern Observatory, D-85748 Garching, Germany 2 Observatoire de Paris-Meudon, F-92125 Meudon Cedex, France

We report on the first high-resolution spectroscopy of 10 giants in LMC Globular Clusters in a wide age range, obtained with the newly commissioned spectrograph UVES at VLT UT2. These observations are used to derive oxygen and iron content of these clusters, and the abundances are then used to cast a more precise view, not only on the age-metallicity relation in the LMC, but also on the chemical evolution of this dwarf irregular galaxy.

Accepted by: Astronomy & Astrophysics For preprints, contact vhill@eso.org

The Optical Gravitational Lensing Experiment. Stellar Distance Indicators in the Magellanic Clouds and Constraints on the Magellanic Cloud Distance Scale

A. Udalski

1 Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland

BVI photometry of the Magellanic Clouds collected during the OGLE-II microlensing experiment makes it possible to study in detail photometric properties of the “major” stellar distance indicators in the Magellanic Clouds. In addition to Cepheids, RR Lyr and red clump stars, which photometry was presented in the earlier OGLE papers, we present the so far most accurate determination of brightness of the tip of the red giant branch in the LMC and SMC.

We analyze the ratios of brightness of all four distance indicators in the LMC and SMC. Additionally, we include in our analysis, when possible, photometric data of the distance indicators in the metal poor Carina dwarf galaxy for which photometry was also collected during the OGLE-II experiment. The analysis is largely differential, free from zero point and extinction uncertainties.

The main conclusion is that the distance scales resulting from all four stellar distance indicators are fully consistent with each other. Therefore the distance scale problem is not a problem resulting from different distances yielded by distance indicators, but rather a problem of the proper zero point of this common distance scale. All four stellar distance indicators have to be treated as an ensemble, and any determination of the zero point for one of them must predict reasonable luminosities of the others. This puts strong constraints on many proposed calibrations of distance indicators. For example, very bright calibrations of Cepheids or RR Lyr can practically be ruled out.

At present, the most likely calibration of the zero point of the common distance scale, which would be consistent with observations of all four distance indicators, is that resulting from faint calibration of RR Lyr stars or calibration of the red clump stars. With the OGLE-II photometry it leads to the distance moduli of (m − M)_{LMC} = 18.24 mag, (m − M)_{SMC} = 18.75 mag and (m − M)_{CAR} = 19.94 mag for the LMC, SMC and Carina dwarf galaxy, respectively. The systematic uncertainty is of the order of 0.07 mag, while the standard deviation of four determinations of only 0.02 mag.

We also analyze the difference of distance moduli between the SMC and LMC. The average value is Δ(m − M)_{SMC−LMC} = 0.50 ± 0.03 mag from four independent measurements. This very good agreement allows us to draw conclusions on the interstellar extinction in the Magellanic Clouds. Our photometric data also provide constraints on the properties of red clump stars.

Finally, the differential comparison of brightness of Cepheids with brightness of other distance indicators in the Magellanic Clouds, and preliminary observations of Cepheids in the IC1613 galaxy indicate no dependence of the zero point of Cepheid Period–Luminosity relation on metallicity.

Accepted by: Acta Astronomica Vol. 50, 279.

For preprints, contact udalski@astrouw.edu.pl
Also available from the URL http://www.arXiv.org/abs/astro-ph/0010151

Abstracts of Non-Refereed Papers

The great circle of the Magellanic Stream

N.V.Bystrova^{1}

1 St.Petersburg Branch of the Special Astrophysical Observatory Russ.Acad.Sci. Pulkovo, SAO, 196140 St.Petersburg Russia

The Zone II on the great circle of the Magellanic Stream was chosen as azimuthally opposed to the Zone I studied and published earlier. The preliminary results about the character of the HI emission in the Zone II are described and some comparisons are made with the Warps of the gaseous disk of our Galaxy.

To appear in: The proceedings of the Conference “Galaxy disks and disk galaxies,” sponsored by the Vatican Observatory. June 12 - 16, 2000, Rome, Italy, at the Pontifical Gregorian University. J. Funes and E. M. Corsini (eds.). To be published by the ASP. For preprints, contact N.V.Bystrova bnv@fsao.spb.su

Rapid Photometry of Supernova 1987A: A 2.14 ms Pulsar?

J. Middleditch^{1}, J. A. Kristian^{2,3}, W. E. Kunkel^{2}, K. M. Hill^{4}, R. D. Watson^{4}, R. Lucinio^{5}, J. N. Imamura^{6}, T. Y. Steiman-Cameron^{7}, A. Shearer^{8}, R. Butler^{8} M. Redfern^{9}, A. C. Danks^{10}

1 Los Alamos National Lab., MS B256, Los Alamos, NM 87545 2 Observatories of the Carnegie Institution of Washington, 813 Santa Barbara St., Pasadena, CA 91101 3 Deceased 4 School of Mathematics and Physics, University of Tasmania, Hobart, Tasmania, Australia

5 Independent Consultant and Division of Geological and Planetary Sciences, 170-25, California Institute of Technology, Pasadena, CA 91125 6 Institute for Theoretical Science, University of Oregon, Eugene, OR 97403 7 NASA/Ames Research Center, MS 245-3, Moffett Field, CA 94035 8 Department of Information Technology, National University of Ireland, Galway, Ireland 9 Dept. of Physics, National University of Ireland, Galway, Ireland 10 RPSC/NASA Goddard Space Flight Center, Building 21, Code 683.0, Greenbelt, MD 20771

We have monitored Supernova 1987A in optical/near-infrared bands from a few weeks following its birth until the present time in order to search for a pulsar remnant. We have found an apparent pattern of emission near the frequency of 467.5 Hz – a 2.14 ms pulsar candidate, first detected in data taken on the remnant at the Las Campanas Observatory (LCO) 2.5-m Dupont telescope during 14-16 Feb. 1992 UT. We detected further signals near the 2.14 ms period on numerous occasions over the next four years in data taken with a variety of telescopes, data systems and detectors, at a number of ground- and space-based observatories. The sequence of detections of this signal from Feb. ‘92 through August ‘93, prior to its apparent subsequent fading, is highly improbable (< 10^{−10} for any noise source). We also find evidence for modulation of the 2.14 ms period with a ∼1,000 s period which, when taken with the high spindown of the source (2-3×10^{−10} Hz/s), is consistent with precession and spindown via gravitational radiation of a neutron star with a non-axisymmetric oblateness of ∼10^{−6}, and an implied gravitational luminosity exceeding that of the Crab Nebula pulsar by an order of magnitude.

To appear in: Proceedings of the conference: Supernova 1987A – Ten Years After For preprints, contact jon@lanl.gov

Editors’ note: The refereed version of this paper was published in New Astronomy, 5, 243 (2000).

Theoretical expectations for clump red giants as distance indicators

Léo Girardi

Dipartimento di Astronomia, Vicolo dell’Osservatorio 5, I-35122 Padova, Italy

Variations of ∼ 0.4 mag are expected in the I-band absolute magnitude of red clump giants, M_{I}^{RC}, as a function of both stellar age and metallicity. This regardless of some potential theoretical uncertainties. Due to the quite large differences in mean ages and metallicities of clump stars among galaxies, systematic changes (also amounting up to ∼ 0.4 mag) come out in their M_{I}^{RC}. These numbers also indicate a distance to the LMC that is not necessarily “short”.

To appear in: Highlights of Astronomy, Vol. 12, proc. of the XXIVth IAU General Assembly, Joint Discussion JD13: HIPPARCOS and the luminosity calibration of the nearer stars, ed. T. Lloyd-Evans

For preprints, contact Lgirardi@pd.astro.it
Also available from the URL http://www.arXiv.org/abs/astro-ph/0010184

Rotation of the Disk of the Large Magellanic Cloud

David R. Alves^{1}

1 Space Telescope Science Institute, USA

The rotation of the disk of the LMC is derived from the radial velocities of 975 carbon stars. The radial component of the disk’s velocity ellipsoid is σ_{R} = 28 ± 14 km s^{−1}. The disk rotates with a true maximum circular velocity of 79 km s^{−1} after correcting for asymmetric drift. For a typical submaximal disk, the LMC’s dark halo has a mass of 1 × 10^{10}M_{⊙} inside of 10 kpc. If the LMC microlensing lens population is a halo of stars in the LMC, its mass-to-light ratio is > 10.

To appear in: Galaxy Disks and Disk Galaxies, eds. J.G. Funes S.J. and E.M. Corsini For preprints, contact alves@stsci.edu

Also available from the URL http://xxx.lanl.gov/astro-ph/0010098

Job Opportunities

TENURE TRACK FACULTY POSITION -

OBSERVATIONAL ASTRONOMY/ASTROPHYSICS

RICE UNIVERSITY

The Department of Physics & Astronomy at Rice University invites applications for a tenure-track faculty position in the field of observational astronomy/astrophysics related to star formation and origins. Specific research areas of interest include protostellar objects and outflows, young clusters and associated emission nebulae and molecular clouds, and global star formation processes on galactic and extragalactic scales. Candidates with multiwavelength research experience (UV–optical–IR–radio) and/or instrumentation experience (instrument development or participation in a development team) are particularly encouraged to apply.

This appointment is expected to be made at the assistant professor level, but outstanding applicants can receive consideration for accelerated promotion or appointment as associate professor with tenure. Review of applications will begin on 1 December 2000 and continue until the position is filled. Interested applicants should send a resume and statement of current research activities and future interests, reprints of two recent publications, and arrange to have three letters of reference sent to the address below.

Rice University is one of the premier small private universities in the nation (www.rice.edu) with an exceptionally talented undergraduate student body and a strong commitment to excellence in research and graduate education. The Department of Physics & Astronomy

                          (dacnet.rice.edu/depts/ricephys/)

currently has several faculty engaged in research in observational and theoretical astrophysics. Rice University encourages applications from qualified women and minorities. AAE/EOE.

Prof. R. Dufour, Astronomy Search Committee

Department of Physics & Astronomy

MS 108 Rice University

Houston, TX 77005-1892

USA

Tel: 713-348-4939

FAX: 713-348-5143

E-mail Inquiries: rjd@rice.edu POSTDOCTORAL RESEARCH ASSISTANT

The UNIVERSITY of MANCHESTER

INSTITUTE of SCIENCE and TECHNOLOGY

The Astrophysics Group within the Department of Physics at UMIST has an opening for a Post-Doctoral Research Assistant in the field of observations of evolved stars, starting as soon as possible, for a period of one year renewable until October 2003.

The Group (http://saturn.phy.umist.ac.uk:8000) includes G.A. Fuller, M.D. Gray, T.J. Millar and A.A. Zijlstra, and is a world leader in the field of molecular astrophysics, including the physics and chemistry of interstellar clouds and circumstellar regions, star formation and evolved stars and their circumstellar envelopes.

We are seeking someone with a strong interest in observations to work with Dr A.A. Zijlstra on the evolution of Asymptotic Giant Branch stars in nearby galaxies, and the formation of circumstellar dusty envelopes around these stars. Applicants with a theoretical background, and those who have observational experience relevant to other areas of our research, will also be considered.

The appointment will be made on the PDRA 1A research scale, starting salary £18185 per annum (under review). Applicants should send a letter outlining their research interests, a CV, and the names and addresses of two referees, to

Dr A.A. Zijlstra

Department of Physics, UMIST PO Box 88 Manchester M60 1QD UK

   E-mail: a.zijlstra@umist.ac.uk

(Tel. +44-(0)161-200-3925)

Deadline for receipt of the application is 1 November 2000. An Equal Opportunities Employer

Meeting Announcement

IAU Symposium 207 Extragalactic Star Clusters

Time and venue: March 12–16, 2001, Pucón, Chile

Scientific rationale: This IAU Symposium aims at exploring the diversity and common properties of star clusters ranging from ancient globulars to young starburst clusters in different environments. We want to bring together researchers working on the many different aspects of extragalactic star clusters in a variety of different wavelength ranges in order to gain a comprehensive picture of star cluster formation and evolution and their role in the evolutionary framework of their parent galaxies. Topics to be addressed include:

• Star cluster formation in its entirety ranging from short-lived, low-density OB associations to open clusters, rich populous clusters, and to very massive, dense super star clusters, the possible progenitors of today’s globular clusters.

• Star cluster formation in all possible environments including low-mass dwarf galaxies, irregular and spiral galaxies, starburst galaxies, interacting and merging galaxies or galaxies that suffered such events on a large scale in the past, isolated ellipticals and ellipticals in galaxy cluster environments.

• The preconditions for cluster formation as a function of environment and external influences, questions of variations of the initial mass function, the question of the existence and length of a period of preferred globular cluster formation, cluster life times and destruction mechanisms.

• The use of star clusters, especially globular clusters, to study galaxy formation and evolution, galaxy dynamics, and the distance scale.

For more details, please check the meeting webpage given below.

SOC: Beatriz Barbuy, Jean Brodie, Gary Da Costa, Bruce Elmegreen, Uta Fritze-von Alvensleben, Juan Forte, Doug Geisler (Co-Chair), Eva Grebel (Co-Chair), Luis Ho, Rob Kennicutt, Myung-Gyoon Lee, Georges Meylan.

Invited Speakers (partial list): Gustavo Bruzual, Roberto Buonanno, Cathy Clarke, Francesca D’Antona, Rosa González-Delgado, Yuri Efremov, Mike Fall, Ken Freeman, Jay Gallagher, Reinhard Genzel, Bill Harris, Jim Hesser, Pascale Jablonka, Markus Kissler-Patig, Elizabeth Lada, Hyung Mok Lee, Mónica Rubio, Ram Sagar, François Schweizer, Alistair Walker, Steve Zepf.

Contact: Eva K. Grebel, Max Planck Institute for Astronomy, starclus@mpia-hd.mpg.de Doug Geisler, Universidad de Concepción, doug@kukita.cfm.udec.cl

Deadline for early registration: Nov 15, 2000 (see website)

URL: http://www.mpia-hd.mpg.de/~starclus/